OpenStudy (anonymous):

ΔABC has three sides with these lengths: AB = 9, BC = 40, and CA = 41. What is the value of cos C?

OpenStudy (anonymous):

 To find the angle measure, the inverse of cosine must be used. Recall from SOH CAH TOA that Cosine is Adjacent/Hypotenuse So our set up is: \[\cos^{-1} \theta=\frac{ 40 }{ 41 }\]